What will be the mass concentration of the HCl solution if 400 cm3 of water is added to 100 cm3 of a solution with a mass concentration of 10 g / dm3?


1
Expert's answer
2021-05-12T06:30:32-0400

(100 cm3) × (1dm3 / 1000 cm3) = 0.1 dm3


Solution:

(0.1 dm3 HCl) × (10 g HCl / 1 dm3 HCl) = 1 g HCl

Mass of solute = 1 g HCl

Volume of solution = 400 cm3 + 100 cm3 = 500 cm3 = 0.5 dm3


Mass concentration = Mass of solute / Volume of solution

Therefore,

Mass concentration = (1 g) / (0.5 dm3) = 2 g/dm3

Mass concentration = 2 g/dm3


Answer: The mass concentration of the HCl solution will be 2 g/dm3